OpenStudy (anonymous):

So I have figured out half this question. Besides For number 4 where would the L go in the final answer if we are solving for D. > > The total leverage of a business is given by the formula T= L + D/1 where T = total leverage, L = liabilities, D = debts, and I = income. A business can use the inequality T > L + D/1 to increase their buying power. Rewrite this formula to solve for D. Show all steps in your work. T = (L+D)/1 T *1 = (L+D)/1*1 1T = L + D -L -L 1T = D

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

In the original question is the (L+D)/I or is it L+(D/I) ? if that makes sense

OpenStudy (anonymous):

It's T= (L+D)/l

OpenStudy (anonymous):

yeah u almost had it then TI = L + D

OpenStudy (anonymous):

then you need to get the L away from the D to leave D on its own

OpenStudy (anonymous):

so D = TI - L

OpenStudy (anonymous):

Wait .. I'm confused now ! Haha Can you write out the full simplified answer ?

OpenStudy (anonymous):

you already had it worked out to IT = L + D

OpenStudy (anonymous):

which was right

OpenStudy (anonymous):

so you just need to get that L on the right hand side of the = sign over to the left hand side

OpenStudy (anonymous):

so you just minus it from the other side

OpenStudy (anonymous):

giving you IT - L = D

OpenStudy (anonymous):

so Is that the final answer

OpenStudy (anonymous):

@alby69

OpenStudy (anonymous):

yeah, i didnt see the second half of that question lol 1 sec

OpenStudy (anonymous):

IT - L = D IT - L > D so you could rewrite it this way because we normally have the letter we are finding on the left, not that it reall matters, but i would write it like this: D< IT - L